OpenStudy (clamin):

PLEASE HELP!! MEDAL!! Multiplying rational expression 2a2bb2c×ba

OpenStudy (clamin):

\[\frac{ 2a ^{2}b }{ b ^{2} c}\times \frac{ b }{ a }\]

OpenStudy (anonymous):

2a/c

OpenStudy (anonymous):

2ab/bc

OpenStudy (clamin):

what are the solutions??

OpenStudy (clamin):

@Schen421 @RogerYE

OpenStudy (anonymous):

just cancel out the same part from denominator and numerator

OpenStudy (anonymous):

\[\frac{ 2ab }{ bc }\]